OpenStudy (anonymous):

yo yo yo help me bro!!!!!What is the solution to this equation? x + 8 + 5x = 32 - 2x A. x = 5 B. x = 6 C. x = 10 D. x = 3

OpenStudy (anonymous):

Simplifying x + 8 + 5x = 32 + -2x Reorder the terms: 8 + x + 5x = 32 + -2x Combine like terms: x + 5x = 6x 8 + 6x = 32 + -2x Solving 8 + 6x = 32 + -2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '2x' to each side of the equation. 8 + 6x + 2x = 32 + -2x + 2x Combine like terms: 6x + 2x = 8x 8 + 8x = 32 + -2x + 2x Combine like terms: -2x + 2x = 0 8 + 8x = 32 + 0 8 + 8x = 32 Add '-8' to each side of the equation. 8 + -8 + 8x = 32 + -8 Combine like terms: 8 + -8 = 0 0 + 8x = 32 + -8 8x = 32 + -8 Combine like terms: 32 + -8 = 24 8x = 24 Divide each side by '8'. x = 3 Simplifying x = 3 D

OpenStudy (anonymous):

Shift all x on one side:)

OpenStudy (anonymous):